# VMamba: Visual State Space Model

### v3

@@ Abstract
Designing computationally efficient network architectures persists as an ongoing necessity in computer vision. In this paper, we transplant Mamba, a state-space language model, into VMamba, a vision backbone that works in linear time complexity. At the core of VMamba lies a stack of Visual State-Space (VSS) blocks with the 2D Selective Scan (SS2D) module. By traversing along four scanning routes, SS2D helps bridge the gap between the ordered nature of 1D selective scan and the non-sequential structure of 2D vision data, which facilitates the gathering of contextual information from various sources and perspectives. Based on the VSS blocks, we develop a family of VMamba architectures and accelerate them through a succession of architectural and implementation enhancements. Extensive experiments showcase VMamba’s promising performance across diverse visual perception tasks, highlighting its advantages in input scaling efficiency compared to existing benchmark models. Source code is available at https://github.com/MzeroMiko/VMamba.
@@ 1 Introduction
Visual representation learning stands as a fundamental research area in computer vision, which has witnessed remarkable progress in the era of deep learning. To represent complex patterns in vision data, two primary categories of backbone networks, i.e., Convolution Neural Networks (CNNs) [49, 28, 30, 54, 38] and Vision Transformers (ViTs) [13, 37, 58, 68], have been proposed and extensively utilized in a variety of visual tasks. Compared to CNNs, ViTs generally demonstrate superior learning capabilities on large-scale data due to the integration of the self-attention mechanism [59, 13]. However, the quadratic complexity of self-attention w.r.t. the number of tokens introduces substantial computational overhead in downstream tasks involving large spatial resolutions.
@@
To tackle this challenge, considerable efforts have been made to enhance the efficiency of attention computation [55, 37, 12]. However, existing approaches either impose limitations on the size of the effective receptive field [37] or experience evident performance degradation across diverse tasks [31, 62]. This motivates us to develop a novel architecture for vision data, preserving the inherent advantages of the vanilla self-attention mechanism, i.e., global receptive fields and dynamic weighting parameters [23].
@@
Recently, Mamba [17], a novel State Space Model (SSM) [17, 44, 61] in the field natural language processing (NLP), has emerged as a highly promising approach for long sequence modeling with linear complexity. Drawing inspiration from this advancement, we introduce VMamba, a vision backbone integrating SSM-based blocks to facilitate efficient visual representation learning. However, the core algorithm of Mamba, i.e., the parallelized selective scan operation, is essentially designed for processing one-dimensional sequential data. This poses a challenge when attempting to adapt it for processing vision data, which inherently lacks a sequential arrangement of visual components. To address this issue, we propose 2D Selective Scan (SS2D), a four-way scanning mechanism tailored for spatial domain traversal. In contrast to the self-attention mechanism (Figure 1 (a)), SS2D ensures that each image patch gains contextual knowledge exclusively through a compressed hidden state computed along the corresponding scanning path (Figure 1 (b)), thereby reducing the computational complexity from quadratic to linear.
@@
Figure 1: Comparison of correlation establishment between image patches via (a) self-attention and (b) the proposed 2D-Selective-Scan (SS2D). Red boxes indicate the query image patch, with patch opacity representing the degree of information loss.
@@
Upon the VSS blocks, we develop a family of VMamba architectures (i.e., VMamba-Tiny/Small/Base) and accelerate them through a series of architectural enhancements and implementation optimizations. Compared to benchmark vision models built on CNN (ConvNeXt [38]), ViT (Swin [37] and HiViT [68]), and SSM (S4ND [45] and Vim [71]), VMamba consistently achieves superior image classification accuracy on ImageNet-1K [9] across model scales. Specifically, VMamba-Base achieves a top-1 accuracy of $83.9\%$, surpassing Swin by $+0.4\%$, with a throughput exceeding that of Swin by a substantial margin over $40\%$ ($646$ vs. $458$). The superiority of VMamba extends across various downstream tasks, with VMamba-Tiny/Small/Base achieving $47.3\%$/$48.7\%$/$49.2\%$ mAP in object detection on COCO [34] (1×1$\times$1 × training schedule). This outperforms Swin by $4.6\%$/$3.9\%$/$2.3\%$ and ConvNeXt by $3.1\%$/$3.3\%$/$2.2\%$, respectively. As for single-scale semantic segmentation on ADE20K [70], VMamba-Tiny/Small/Base achieves $47.9\%$/$50.6\%$/$51.0\%$ mIoU, which surpasses Swin by $3.4\%$/$3.0\%$/$2.9\%$ and ConvNeXt by $1.9\%$/$1.9\%$/$1.9\%$, respectively. Furthermore, unlike ViT-based models, which experience quadratic growth in computational complexity with the number of input tokens, VMamba exhibits linear growth in FLOPs while maintaining comparable performance. This underscores its state-of-the-art input scalability.
@@
  • We propose VMamba, an SSM-based vision backbone network for visual representation learning with linear time complexity. A series of improvements in architectural design and implementation details are adopted to improve the inference speed of VMamba.
  • We introduce 2D Selective Scan (SS2D) to bridge the gap between 1D array scanning and 2D plane traversal, facilitating the extension of selective SSM to process vision data.
  • Without bells and whistles, VMamba demonstrates promising performance across a range of visual tasks, including image classification, object detection, and semantic segmentation. It also exhibits remarkable adaptability w.r.t. the length of the input sequence, showcasing linear growth in computational complexity.
@@ 2 Related Work
Starting from AlexNet [32], significant efforts have been dedicated to enhancing the modeling capability [49, 53, 28, 30] and computational efficiency [29, 54, 66, 47] of CNN-based models across a spectrum of visual tasks. More sophisticated operators, such as the depth-wise convolution [29] and deformable convolution [5, 72], have been proposed for increasing the flexibility and efficacy of CNNs. Recently, drawing inspiration from the success of Transformers [60], modern CNNs [38] have demonstrated promising performance by incorporating long-range dependencies [11, 48, 35] and dynamic weights [24] into their architectural design.
@@
As the most representative pioneering work, ViT [13] explores the effectiveness of vision models based on vanilla Transformer architecture and reveals the significance of large-scale pre-training in improving the performance in image classification. To mitigate ViT’s reliance on extremely large datasets, DeiT [58] introduces a teacher-student distillation strategy to transfer knowledge encapsulated in CNN models to ViTs, showcasing the importance of the inductive bias in visual perception. Following this line of thought, subsequent studies propose hierarchical ViTs [37, 12, 63, 40, 68, 56, 6, 10, 69, 1].
Another line of research focuses specifically on enhancing the computational efficiency of the self-attention mechanism, which serves as the cornerstone of ViTs. By formulating self-attention as the linear dot-product of kernel feature maps, Linear Attention [31] leverages the associativity property of matrix products to reduce the computational complexity from quadratic to linear. GLA [67] further introduces a hardware-efficient variant of linear attention, which balances memory movement against parallelizability. RWKV [46] also leverages the linear attention mechanism to combine the parallelizable training of transformers with the efficient inference of recurrent neural networks (RNNs). RetNet [52] integrates an additional gating mechanism to enable a parallelizable computation path as an alternative to recurrence, and RMT [15] extends it for visual representation learning by incorporating the temporal decay mechanism into the spatial domain.
@@
Despite the widespread adoption of ViT architectures in vision tasks, the quadratic complexity of self-attention presents a significant challenge when dealing with long input sequences, (e.g., high-resolution images). Among the attempts to improve the scaling efficiency [8, 7, 46, 52, 42], SSMs have emerged as compelling alternatives to Transformers, attracting significant attention from the community. Gu et al. [21] demonstrated the potential of SSM-based models in handling the long-range dependency with the HiPPO initialization [18]. To improve the practical feasibility, S4 [20] proposed to normalize the parameter matrices into a diagonal structure. Subsequently, various structured SSM models have emerged, each introducing different architectural enhancements. These include models with complex-diagonal structures [22, 19], support for multiple-input multiple-output [50], decomposition of diagonal plus low-rank operations [25], and selection mechanisms [17]. These advancements have also been integrated into larger representation models [44, 42, 16], further demonstrating the versatility and scalability of structured state space models in various applications. While these models primarily focus on applying SSMs to long-range and sequential data like text and speech, there has been limited exploration of SSMs’ application to vision data with two-dimensional structures.
@@ 3 Preliminaries
Originating from the Kalman filter [33], SSMs can be regarded as linear time-invariant (LTI) systems that map the input stimulation $u(t)\in\mathbb{R}$ to response $y(t)\in\mathbb{R}$ through the hidden state $\mathbf{h}(t)\in\mathbb{R}^{N}$. Concretely, continuous-time SSMs can be formulated as linear ordinary differential equations (ODEs) as follows,
@@
To be integrated into deep models, continuous-time SSMs must undergo discretization in advance. Concretely, consider the time interval $[t_{a},t_{b}],$the analytic solution to the hidden state variable $\mathbf{h}(t)|_{t={t_{b}}}$ can be expressed as
@@
To tackle the limitation of LTI SSMs (Eq. 1) in capturing the contextual information, Gu et al. [17] propose a novel parameterization method for SSMs that integrates an input-dependent selection mechanism (referred to as S6). However, in the case of selective SSMs, the time-varying weighting parameters present a challenge for efficient computation of hidden states, as convolutions do not accommodate dynamic weights and are consequently rendered inapplicable. Nevertheless, as the recurrence relation of $h_{b}$ in Eq. 3 can be derived, the response $y_{b}$ can still be efficiently computed using associative scan algorithms [2, 43, 51] with linear complexity (the detailed explanation is deferred to Appendix B).
@@ 4.1 Network Architecture
We develop VMamba at three scales: VMamba-Tiny, VMamba-Small, and VMamba-Base (referred to as VMamba-T, VMamba-S, and VMamba-B, respectively). An overview of the architecture of VMamba-T is illustrated in Figure 3 (a), and detailed configurations are provided in Appendix  E. The input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ is first partitioned into patches by a stem module, resulting in a 2D feature map with the spatial dimension of $H/4\times W/4$. Subsequently, multiple network stages are employed to create hierarchical representations with resolutions of $H/8\times W/8$, $H/16\times W/16$, and $H/32\times W/32$. Each stage comprises a down-sampling layer (except for the first stage) followed by a stack of Visual State Space (VSS) blocks.
The VSS blocks serve as the visual counterpart to Mamba blocks [17] (Figure 3 (b)) for representation learning. The initial architecture of VSS blocks (referred to as the ‘vanilla VSS Block’ in Figure 3 (c)) is formulated by substituting the S6 module, which stands as the core of Mamba in concurrently achieving global receptive fields, dynamic weights (i.e., selectivity), and linear complexity, with the newly proposed 2D-Selective-Scan (SS2D) module (to be introduced in the following subsection). To further enhance computational efficiency, we eliminate the entire multiplicative branch (encircled by the red box in Figure 3 (c)), as the effect of the gating mechanism has been achieved by the selectivity of SS2D. Consequently, the resulting VSS block (depicted in Figure 3 (d)) consists of a single network branch with two residual modules, mimicking the architecture of a vanilla Transformer block [60]. Throughout this paper, all results are obtained using VMamba models built with VSS blocks in this architecture.
Figure 2: Illustration of 2D-Selective-Scan (SS2D). Input patches are traversed along four different scanning paths (Cross-Scan), and each sequence is independently processed by distinct S6 blocks. Subsequently, the results are merged to construct a 2D feature map as the final output (Cross-Merge).
Figure 3: Left: Illustration of (a) the overall architecture of VMamba, and (b) - (d) the structure of Mamba and VSS blocks. Right: Comparison of variants of VMamba and benchmark methods in classification accuracy and computational efficiency.
@@
While the sequential nature of the scanning operation in S6 aligns well with NLP tasks involving temporal data, it poses a significant challenge when applied to vision data, which is inherently non-sequential and encompasses spatial information (e.g., local texture and global structure). To address this issue, S4ND [45] reformulates SSM with convolutional operations, directly extending the kernel from 1D to 2D through outer-product. However, such modification prevents the weights from being input-independent, resulting in a limited capacity for capturing contextual information. Therefore, we adhere to the selective scan approach [17] for input processing, and propose the 2D-Selective-Scan (SS2D) module to adapt S6 to vision data without compromising its advantages.
As illustrated in Figure 2, data forwarding in SS2D involves three steps: cross-scan, selective scanning with S6 blocks, and cross-merge. Given the input data, SS2D first unfolds input patches into sequences along four distinct traversal paths (i.e., Cross-Scan), processes each patch sequence using a separate S6 block in parallel, and subsequently reshapes and merges the resultant sequences to form the output map (i.e., Cross-Merge). By adopting complementary 1D traversal paths, SS2D enables each pixel in the image to effectively integrate information from all other pixels in different directions, thereby facilitating the establishment of global receptive fields in the 2D space.
@@ 4.3 Accelerating VMamba
As shown in Figure 3 (e), the VMamba-T model with vanilla VSS blocks (referred to as ‘Vanilla VMamba’) achieves a throughput of $426$ images/s and consists of $22.9$M parameters with $5.6$G FLOPs. Despite the state-of-the-art $82.2\%$ classification accuracy (in the tiny-level, outperforms Swin-T [37] by $0.9\%$), the low throughput and high memory overhead pose significant challenges for the practical deployment of VMamba.

### v4

@@ Abstract
Designing computationally efficient network architectures remains an ongoing necessity in computer vision. In this paper, we adapt Mamba, a state-space language model, into VMamba, a vision backbone with linear time complexity. At the core of VMamba is a stack of Visual State-Space (VSS) blocks with the 2D Selective Scan (SS2D) module. By traversing along four scanning routes, SS2D bridges the gap between the ordered nature of 1D selective scan and the non-sequential structure of 2D vision data, which facilitates the collection of contextual information from various sources and perspectives. Based on the VSS blocks, we develop a family of VMamba architectures and accelerate them through a succession of architectural and implementation enhancements. Extensive experiments demonstrate VMamba’s promising performance across diverse visual perception tasks, highlighting its superior input scaling efficiency compared to existing benchmark models. Source code is available at https://github.com/MzeroMiko/VMamba
@@ 1 Introduction
Visual representation learning remains as a fundamental research area in computer vision that has witnessed remarkable progress in the era of deep learning. To represent complex patterns in vision data, two primary categories of backbone networks, i.e., Convolutional Neural Networks (CNNs) [49, 27, 29, 53, 37] and Vision Transformers (ViTs) [13, 36, 57, 66], have been proposed and extensively utilized in a variety of visual tasks. Compared to CNNs, ViTs generally demonstrate superior learning capabilities on large-scale data due to their integration of the self-attention mechanism [58, 13]. However, the quadratic complexity of self-attention w.r.t. the number of tokens imposes substantial computational overhead in downstream tasks involving large spatial resolutions.
@@
To address this challenge, significant efforts have been made to improve the efficiency of attention computation [54, 36, 12]. However, existing approaches either restrict the size of the effective receptive field [36] or suffer from notable performance degradation across various tasks [30, 60]. This motivates us to develop a novel architecture for vision data, while maintaining the inherent advantages of the vanilla self-attention mechanism, i.e., global receptive fields and dynamic weighting parameters [23].
@@
Recently, Mamba [17], a innovative State Space Model (SSM) [17, 43, 59, 71, 48], in the field of natural language processing (NLP), has emerged as a promising approach for long-sequence modeling with linear complexity. Drawing inspiration from this advancement, we introduce VMamba, a vision backbone that integrates SSM-based blocks to enable efficient visual representation learning. However, the core algorithm of Mamba, i.e., the parallelized selective scan operation, is essentially designed for processing one-dimensional sequential data. This presents a challenge when adapting it for processing vision data, which lacks an inherent sequential arrangement of visual components. To address this issue, we propose 2D Selective Scan (SS2D), a four-way scanning mechanism designed for spatial domain traversal. In contrast to the self-attention mechanism (Figure 1 (a)), SS2D ensures that each image patch acquires contextual knowledge exclusively through a compressed hidden state computed along its corresponding scanning path (Figure 1 (b)), thereby reducing the computational complexity from quadratic to linear.
@@
Figure 1: Comparison of the establishment of correlations between image patches through (a) self-attention and (b) the proposed 2D-Selective-Scan (SS2D). The red boxes indicate the query image patch, with its opacity representing the degree of information loss.
@@
Building on the VSS blocks, we develop a family of VMamba architectures (i.e., VMamba-Tiny/Small/Base) and enhance their performance through architectural improvements and implementation optimizations. Compared to benchmark vision models built on CNNs (ConvNeXt [37]), ViTs (Swin [36], HiViT [66]), and SSMs (S4ND [44], Vim [69]), VMamba consistently achieves higher image classification accuracy on ImageNet-1K [9] across various model scales. Specifically, VMamba-Base achieves a top-1 accuracy of $83.9\%$, surpassing Swin by $+0.4\%$, with a throughput exceeding Swin’s by a substantial margin over $40\%$ ($646$ vs. $458$). VMamba’s superiority extends across multiple downstream tasks, with VMamba-Tiny/Small/Base achieving $47.3\%$/$48.7\%$/$49.2\%$ mAP in object detection on COCO [33] (1×1$\times$1 × training schedule). This outperforms Swin by $4.6\%$/$3.9\%$/$2.3\%$ and ConvNeXt by $3.1\%$/$3.3\%$/$2.2\%$, respectively. As for single-scale semantic segmentation on ADE20K [68], VMamba-Tiny/Small/Base achieves $47.9\%$/$50.6\%$/$51.0\%$ mIoU, which surpasses Swin by $3.4\%$/$3.0\%$/$2.9\%$ and ConvNeXt by $1.9\%$/$1.9\%$/$1.9\%$, respectively. Furthermore, unlike ViT-based models, which experience quadratic growth in computational complexity with the number of input tokens, VMamba exhibits linear growth in FLOPs while maintaining comparable performance. This demonstrates its state-of-the-art input scalability.
@@
  • We propose VMamba, an SSM-based vision backbone for visual representation learning with linear time complexity. A series of architectural and implementation improvements are adopted to enhance the inference speed of VMamba.
  • We introduce 2D Selective Scan (SS2D) to bridge 1D array scanning and 2D plane traversal, enabling the extension of selective SSMs to process vision data.
  • VMamba achieves promising performance across various visual tasks, including image classification, object detection, and semantic segmentation. It also exhibits remarkable adaptability w.r.t. the length of the input sequence, showcasing linear growth in computational complexity.
@@ 2 Related Work
Since AlexNet [31], considerable efforts have been devoted to enhancing the modeling capabilities [49, 52, 27, 29] and computational efficiency [28, 53, 64, 46] of CNN-based models across various visual tasks. Sophisticated operators like depth-wise convolution [28] and deformable convolution [5, 70] have been introduced to increase the flexibility and efficacy of CNNs. Recently, inspired by the success of Transformers [58], modern CNNs [37] have shown promising performance by integrating long-range dependencies [11, 47, 34] and dynamic weights [23] into their designs.
@@
As a pioneering work, ViT [13] explores the effectiveness of vision models based on vanilla Transformer architecture, highlighting the importance of large-scale pre-training for image classification performance. To reduce ViT’s dependence on large datasets, DeiT [57] introduces a teacher-student distillation strategy, transferring knowledge from CNNs to ViTs and emphasizing the importance of inductive bias in visual perception. Following this approach, subsequent studies propose hierarchical ViTs [36, 12, 61, 39, 66, 55, 6, 10, 67, 1].
Another research direction focuses on improving the computational efficiency of self-attention, which serves as the cornerstone of ViTs. Linear Attention [30] reformulates self-attention as a linear dot-product of kernel feature maps, using the associativity property of matrix products to reduce computational complexity from quadratic to linear. GLA [65] introduces a hardware-efficient variant of linear attention that balances memory movement with parallelizability. RWKV [45] also leverages the linear attention mechanism to combine parallelizable transformer training with the efficient inference of recurrent neural networks (RNNs). RetNet [51] adds a gating mechanism to enable a parallelizable computation path, offering an alternative to recurrence. RMT [15] further extends this for visual representation learning by applying the temporal decay mechanism to the spatial domain.
@@
Despite their widespread adoption in vision tasks, ViT architectures face significant challenges due to the quadratic complexity of self-attention, especially when handling long input sequences (e.g., high-resolution images). In efforts to improve scaling efficiency [8, 7, 45, 51, 41], SSMs have emerged as compelling alternatives to Transformers, attracting significant research attention. Gu et al. [21] demonstrate the potential of SSM-based models in handling the long-range dependencies using the HiPPO initialization [18]. To improve practical feasibility, S4 [20] proposes normalizing the parameter matrices into a diagonal structure. Various structured SSM models have since emerged, each offering distinct architectural enhancements, such as complex-diagonal structures [22, 19], support for multiple-input multiple-output [50], diagonal plus low-rank decomposition [24], and selection mechanisms [17]. These advancements have also been integrated into larger representation models [43, 41, 16], further highlighting the versatility and scalability of structured state space models in various applications. While these models primarily target long-range and sequential data such as text and speech, limited research has explored applying SSMs to vision data with two-dimensional structures.
@@ 3 Preliminaries
Originating from the Kalman filter [32], SSMs are linear time-invariant (LTI) systems that map the input signal $u(t)\in\mathbb{R}$ to the output response $y(t)\in\mathbb{R}$ via the hidden state $\mathbf{h}(t)\in\mathbb{R}^{N}$. Specifically, continuous-time SSMs can be expressed as linear ordinary differential equations (ODEs) as follows,
@@
To be integrated into deep models, continuous-time SSMs must undergo discretization in advance. Concretely, for the time interval $[t_{a},t_{b}]$, the analytic solution of the hidden state variable $\mathbf{h}(t)$ at $t=t_{b}$ can be expressed as
@@
To address the limitation of LTI SSMs (Eq. 1) in capturing the contextual information, Gu et al. [17] propose a novel parameterization method for SSMs, which incorporates an input-dependent selection mechanism (referred to as S6). However, for selective SSMs, the time-varying weighting parameters pose a challenge for efficient computation of hidden states, as convolutions cannot accommodate dynamic weights, making them inapplicable. Nevertheless, since the recurrence relation of $h_{b}$ in Eq. 3 can be derived, the response $y_{b}$ can still be efficiently computed using associative scan algorithms [2, 42, 50], which has linear complexity (see Appendix B for a detailed explanation).
@@ 4.1 Network Architecture
We develop VMamba in three scales: Tiny, Small, and Base (referred to as VMamba-T, VMamba-S, and VMamba-B, respectively). An overview of the architecture of VMamba-T is illustrated in Figure 3 (a), and detailed configurations are provided in Appendix  E. The input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ is first partitioned into patches by a stem module, resulting in a 2D feature map with spatial dimension of $H/4\times W/4$. Without incorporating additional positional embeddings, multiple network stages are employed to create hierarchical representations with resolutions of $H/8\times W/8$, $H/16\times W/16$, and $H/32\times W/32$. Specifically, each stage comprises a down-sampling layer (except for the first stage), followed by a stack of Visual State Space (VSS) blocks.
The VSS blocks serve as the visual counterparts to Mamba blocks [17] (Figure 3 (b)) for representation learning. The initial architecture of VSS blocks (referred to as the ‘vanilla VSS Block’ in Figure 3 (c)) is formulated by replacing the S6 module. S6 is the core of Mamba and achieves global receptive fields, dynamic weights (i.e., selectivity), and linear complexity. We substitute it with the newly proposed 2D-Selective-Scan (SS2D) module, and more details will be introduced in the following subsection. To further enhance computational efficiency, we remove the entire multiplicative branch (highlighted by the red box in Figure 3 (c)), as the effect of the gating mechanism has already been achieved by the selectivity of SS2D. As a result, the improved VSS block (shown in Figure 3 (d)) consists of a single network branch with two residual modules, mimicking the architecture of a vanilla Transformer block [58]. All results in this paper are obtained using VMamba models built with VSS blocks in this architecture.
Figure 2: Illustration of 2D-Selective-Scan (SS2D). Input patches are traversed along four different scanning paths (Cross-Scan), with each sequence independently processed by separate S6 blocks. The results are then merged to construct a 2D feature map as the final output (Cross-Merge).
Figure 3: Left: Illustration of (a) the overall architecture of VMamba, and (b) - (d) the structure of Mamba and VSS blocks. Right: Comparison of VMamba variants and benchmark methods in terms of classification accuracy and computational efficiency.
@@
While the sequential nature of the scanning operation in S6 aligns well with NLP tasks involving temporal data, it poses a significant challenge when applied to vision data, which is inherently non-sequential and encompasses spatial information (e.g., local texture and global structure). To address this issue, S4ND [44] reformulates SSM with convolutional operations, directly extending the kernel from 1D to 2D through the outer-product. However, such modification restricts the weights from being input-dependent, resulting in a limited capacity for capturing contextual information. Therefore, we adhere to the selective scan approach [17] for input processing and propose the 2D-Selective-Scan (SS2D) module to adapt S6 to vision data without compromising its advantages.
Figure 2 illustrates that data forwarding in SS2D consists of three steps: cross-scan, selective scanning with S6 blocks, and cross-merge. Specifically, SS2D first unfolds the input patches into sequences along four distinct traversal paths (i.e., Cross-Scan). Each patch sequence is then processed in parallel using a separate S6 block, and the resultant sequences are reshaped and merged to form the output map (i.e., Cross-Merge). Through the use of complementary 1D traversal paths, SS2D allows each pixel in the image to integrate information from all other pixels across different directions. This integration facilitates the establishment of global receptive fields in the 2D space.
@@ 4.3 Accelerating VMamba
As shown in Figure 3 (e), the VMamba-T model with vanilla VSS blocks (referred to as ‘Vanilla VMamba’) achieves a throughput of $426$ images/s and contains $22.9$M parameters with $5.6$G FLOPs. Despite achieving a state-of-the-art classification accuracy of 82.2% (outperforming Swin-T [36] by 0.9% at the tiny level), the low throughput and high memory overhead present significant challenges for the practical deployment of VMamba.